OpenStudy (anonymous):

determine the factors of x^2 - 8x - 12 ?

OpenStudy (callisto):

First, you can consider the factors of - 12 -12 1 -6 2 -4 3 -3 4 -2 6 -1 12 Then add the factors together and see which gives you -8 For example, -12 + 1 = -11 -1 + 12= 11

OpenStudy (anonymous):

(x+2) (x-6)

OpenStudy (anonymous):

(x+2)(x-6)

OpenStudy (inkyvoyd):

Eh, @Callisto is the only CoC reader

OpenStudy (anonymous):

thank you!!!

OpenStudy (inkyvoyd):

@cassandre , the first one to give the answer like that is an answer sniper. NOt only is that against the rules, but you don't learn anything.